# Efficient Feature Fusion for UAV Object Detection

### v1

@@ IV-D1 Comparisons on Visdrone2019
Models $\mathbf{mAP_{50}}$ $\mathbf{mAP_{50-90}}$ Parameters (M) FLOPs (G)
@@
TABLE III: Detection performance (%) on DOTAv1.5 dataset.
Models $\mathbf{mAP_{50}}$ $\mathbf{mAP_{50-90}}$ Parameters (M)
@@ IV-E Ablation Studies and Analysis
In this section, we use Visdrone2019 as the benchmark dataset and YOLOv10n as the baseline model. Our module is integrated into the baseline model while reducing the number of network layers in the baseline model to ensure that the final parameter count of all methods in the experiment is approximately consistent. This allows us to observe the effects of each module and the issue of imbalance in classification and localization, as discussed in this paper.
In TABLE IV, the $mAP_{50}$ of the YOLOv10n baseline model is 37.3%. The first method, which uses only the FMSA module, upsamples the feature maps before the minimum and medium resolution detection heads using nearest neighbor upsampling and then fuses them with the feature maps before the maximum resolution detection head. As a result, the performance of this single-detection-head network outperforms the three-detection-head YOLOv10n network. This suggests that for a detection head with the highest resolution, there is actually a relative lack of semantic information provided by deeper layers. After supplementing this missing information, the performance of a single detection head is comparable to that of three detection heads.
@@
The third method adds the FDS module on top of the first method, resulting in a significant performance boost. On the one hand, it demonstrates that introducing shallow layer features can significantly enhance the performance of the network. On the other hand, this also supports the problem presented in this paper that the imbalance of classification and localization information exists. As the network architecture becomes deeper, the enhancement of semantic information is often accompanied by a decline in the representation of spatial information. Compared to the second method, in small target detection tasks, the benefit of spatial information in shallow layers slightly outweighs that of semantic information in deep layers.

### v3

@@ IV-D1 Comparisons on VisDrone2019
Model $\mathbf{mAP_{50}}$ $\mathbf{mAP_{50-90}}$ Parameters (M) FLOPs (G)
@@
TABLE III: Detection performance (%) on DOTA-v1.5 dataset.
Model $\mathbf{mAP_{50}}$ $\mathbf{mAP_{50-90}}$ Parameters (M)
@@ IV-E Ablation Studies and Analysis
In this section, we use VisDrone2019 as the benchmark dataset and YOLOv10n as the baseline model. To ensure that the total number of parameters remains approximately the same across all experimental methods, we integrate our module into the baseline model while simultaneously reducing the number of its network layers. This allows us to observe the effects of each module and the issue of imbalance in classification and localization, as discussed in this paper.
In TABLE IV, the $mAP_{50}$ of the YOLOv10n baseline model is 37.3%. The first method upsamples the feature maps using nearest neighbor upsampling before the minimum and medium resolution detection heads. The output is then fused with the feature maps before the maximum resolution detection head via the FMSA module. As a result, the performance of this single-detection-head network outperforms the three-detection-head YOLOv10n network. This suggests that for a detection head with the highest resolution, there is actually a relative lack of semantic information provided by deeper layers. After supplementing this missing information, the performance of a single detection head is comparable to that of three detection heads.
@@
The third method adds the FDS module on top of the first method, resulting in a significant performance boost. On the one hand, it demonstrates that introducing shallow layer features can significantly enhance the performance of the network. On the other hand, this also reveals the problem presented in this paper that the imbalance of classification and localization information exists. As the network architecture becomes deeper, the enhancement of semantic information is often accompanied by a decline in the representation of spatial information. Compared to the second method, in small target detection tasks, the benefit of spatial information in shallow layers slightly outweighs that of semantic information in deep layers.
@@
IV-F Visualization
Figure 7: Comparison of detection results on the VisDrone2019 dataset.
Finally, in Fig. 7 we present representative visualizations from the VisDrone2019 dataset to provide an intuitive demonstration of the impact of our proposed method. These visualizations illustrate the model’s enhanced detection capabilities, offering clear comparative evidence of its performance improvement over the baseline model.
The results demonstrate that our proposed approach achieves significant enhancement in detecting overlapping and extremely small targets across diverse scenarios compared to baseline model. This substantiates the method’s improved precision and effectiveness in complex detection environments, particularly in addressing the challenges introduced by small and occluded objects.
@@ V Conclusion
Acknowledgement
This work is supported by the Sci-Tech Innovation Initiative by the Science and Technology Commission of Shanghai Municipality (24ZR1419000), and the National Science Foundation of China (12471501).